What is 16 - 9u[sup]2[/sup] and the problem is factoring the difference of squares
1
Expert's answer
2011-03-29T11:59:24-0400
As a2-b2 = (a-b)(a+b), 16- 9u2 = 42 - (3u)2 = (4-3u)(4+3u).